Consider taking samples of size 100 from a population with proportion 0.33.

Is the sample size large enough for the Central Limit Theorem to apply so that the sample proportions follow a normal distribution?

a) Yes, np and n(1-p) both >=10.

b) No, np and n(1-p) both >=10.

c) Yes, np and n(1-p) both >=100.

d) No, 100 is never large enough.
Answer #1

Solution :

n = 100

p = 0.33

1 - p = 0.67

np = 100 * 0.33 = 33 and n(1 - p) = 100 * 0.67 = 67 10

a) Yes np and n(1-p) both >=10